OpenStudy (anonymous):

how many eggs are there in 31/2 dozen eggs?

OpenStudy (anonymous):

Okay, so remember that there is twelve in a dozen. This can be written as: dozen = 12 and as you are looking for a number of eggs you can say 31/2 dozen = number of eggs 31/2 * 12 = number of eggs 372/2 = number of eggs number of eggs = 186

OpenStudy (anonymous):

\(\Huge\color{red}\ast\) \(\Huge\mathcal{Welcome\ to\ Openstudy!}\) \(\Huge\color{red}\ast\)